A saleswoman at a car dealership earns $1200 per month plus a 9% commission on her total sales. If she earned $7320 last month, what was the amount of her sales?
Answer by checkley77(12569)

1,200+.09x=7,320
.09x=7,320-1,200
.09x=6,120
x=6,120/.09
x=68,000 is the amount of the sales.
proof
1,200+.09*68,000=7,320
1,200+6,120=7,320
7,320=7,320